OpenStudy (anonymous):

write a recursive definition for the sequence 14, 10, 6, 2.

OpenStudy (anonymous):

What is the continuing difference in the series from term to term?

OpenStudy (anonymous):

Are you asking me or ...

OpenStudy (anonymous):

Yes. The first thing to do is to discover the pattern in the sequence, then you can write an equation that shows how to get the nth term from the one before it.

OpenStudy (anonymous):

For example: 2, 4, 6, 8, ... a(1)=first term = 2 a(2) = a(1)+ 2 =4 a(n) = a(n-1) + 2

OpenStudy (anonymous):

well.... A. a1=14, an-1+4 B.a1=2,an-1-4 c. a1=2, an=an-1+4 D.a1=14. an=an-1-4

OpenStudy (anonymous):

a1 = 14, so you can rule out b and c

OpenStudy (anonymous):

Gone?